# GraphGDel: Constructing and Learning Graph Representations of Genome-Scale Metabolic Models for Growth-Coupled Gene Deletion Prediction

### v6

@@ III-A Constructing Graph Representations of Metabolic Models
As a result, the proposed graph construction pipeline produces a representation that preserves key functional structures while minimizing the confounding influence of topological hubs. The pseudocode of the proposed graph construction pipeline is presented below:
@@
Input: $\mathcal{C}=\{\textit{M},\textit{R},\textit{S},\textit{G},\textit{F}\}$ where:
  • $\textit{M}=\{m_{1},\ldots,m_{|\textit{M}|}\}$: metabolites (including $m_{\text{target}}$)
  • $\textit{R}=\{r_{1},\ldots,r_{|\textit{R}|}\}$: reactions
[…]
  • $\textit{G}=\{g_{1},\ldots,g_{|\textit{G}|}\}$: genes
  • $\textit{F}=\{f_{1},\ldots,f_{|\textit{R}|}\}$: GPR Boolean functions mapping subsets of G to reactions
Currency Metabolite Set:
[…]
@@ III-C1 Metabolite Representation Learning
The Meta-M module is designed to transform raw character strings input, i.e., the SMILES (Simplified Molecular Input Line Entry System) sequences of metabolites into dense, informative feature vectors [smiles]. To effectively capture the complex patterns within these chemical representations, we employ a sequence-to-sequence autoencoder architecture based on Long Short-Term Memory (LSTM) networks [lstm_original, seq2seq]. We selected LSTM as the default model because our goal is to propose a general learning framework for this task and to validate its effectiveness. The LSTM meets these requirements: it is straightforward to implement and provides a strong baseline for future comparisons with more complex replacements, such as Transformer-based models [Transformer].
@@ III-D2 Data Partitioning Strategy
We partitioned the dataset into 90% for training/validation and 10% as an independent test set for final evaluation. Within the training/validation portion, we applied 10-fold stratified splitting. In each fold, approximately 90% of metabolites were used for training and 10% for validation. After selecting the best configuration from the hyperparameter search (see Section III-D3), we retrained the model on the entire training/validation set and evaluated it once on the independent test set. All reported performance metrics are given as the mean and standard deviation across cross-validation folds.
@@ III-D3 Hyperparameters
Hyperparameters were tuned via grid search on the training/validation data using 10-fold stratified cross-validation, repeated five times with different random seeds, to identify the optimal configuration, as summarized in Table IV. The configuration with the highest average validation performance across folds was selected, and the corresponding values were used as the default settings of the proposed framework.
TABLE IV: Grid search for hyperparameter selection. The optimal values are highlighted in underline.
Hyperparameter Search Space
[…]
LSTM hidden dimension {16, 32, 64, 128}
[…]
GCN hidden channels {32, 64, 128, 256}
[…]
Learning rate {0.0001, 0.001, 0.01}
Batch size {4, 8, 16, 32}
Dropout rate {0.1, 0.2, 0.3}
Positional embedding dimension {8, 16, 32}
III-D4 Computational Complexity and Runtime
The computational complexity of the GraphGDel framework is dominated by three components: LSTM autoencoders for sequence processing, GCNs for graph learning, and the combined prediction module.
For LSTM autoencoders in Meta-M and Gene-M modules, the complexity is $\mathcal{O}(L\cdot d\cdot D)$ per layer, where $L$ is sequence length, $d$ is embedding dimension, and $D$ is hidden dimension. The Graph-M module has complexity $\mathcal{O}(|E|\cdot D)$ per GCN layer for a graph with $|E|$ edges. The overall computational complexity for batch size $B$ is:
$\mathcal{O}(B\cdot(L_{g}+L_{m})\cdot d\cdot D+B\cdot|E|\cdot D+B\cdot 3D).$ (25)
Empirically, the framework completes 100 epochs in approximately 10 hours under our experimental conditions, with training time varying based on metabolic model size.
@@ IV-A Overview
  • e_coli_core [orth2010reconstruction] is a simplified metabolic model of Escherichia coli (E. coli) that captures only the essential pathways necessary for core metabolic functions. It is widely used for foundational studies of bacterial metabolism and gene interactions.
  • iMM904 [mo2009connecting] is a more comprehensive model of Saccharomyces cerevisiae (yeast), incorporating a larger set of reactions and metabolites. It is commonly used for analyzing complex eukaryotic metabolic networks.
  • iML1515 [monk2017ml1515] is one of the most detailed metabolic models of E. coli, covering a vast array of metabolic reactions and pathways. It provides a more intricate network compared to e_coli_core and is widely used as a genome-scale model for simulations.
@@
The detailed breakdown of genes, metabolites, and reactions in the studied models is summarized in Table V.
TABLE V: Summary of constraint-based metabolic models used in the computational experiments.
@@ IV-B1 Gene Deletion Strategy Data
Data Source. The gene deletion strategy data used in this study are obtained from the MetNetComp database [tamura2023metnetcomp]. MetNetComp curates systematically simulated results derived from constraint-based models and computes growth-coupled gene deletion strategies that are minimal or maximal with respect to the number of deleted genes.
@@ IV-B2 SMILES Data
Description. The SMILES data represent the metabolites in the studied metabolic models. Each SMILES string encodes the chemical structure of a metabolite, including atoms, bonds, and stereochemistry, as a textual sequence of characters and notations. SMILES is widely used for analyzing chemical properties, such as molecular similarity, substructure searches, and cheminformatics-based predictions [smiles_1, smiles_2, smiles_3].
Data Source. The SMILES data used in this study is obtained from the MetaNetX database [MetaNetX]. MetaNetX is an online platform for accessing, analyzing, and manipulating genome-scale metabolic networks and biochemical pathways. It curates information from public repositories to integrate data on metabolites and related reactions across many species.
For all three metabolic models under study, we download SMILES data for metabolites corresponding to those with available gene deletion strategy data. For metabolites lacking corresponding SMILES entries in the database, we use dummy SMILES sequences as placeholders.
@@ IV-B3 Amino Acid Sequence Data
Description. In metabolic models, gene functions encoded by the GPR rules are ultimately carried out by proteins. Accordingly, we use amino acid (AA) sequences to represent gene information. Each AA sequence corresponds to a protein in the metabolic model and is represented as a linear sequence of amino acids using the standard 20 one-letter amino acid codes. AA sequence data have been widely used to study protein structure and function, including functional annotation, structural prediction, and protein-protein interaction analysis [AA_1, AA_2, AA_3].
Data Source. The AA sequence data used in this study is obtained from the KEGG database [KEGG]. KEGG provides comprehensive information on protein-coding genes, their products, and associated metabolic pathways. The AA sequence data in KEGG is curated from publicly available genomic resources and is continuously updated to ensure accuracy and relevance.
We download AA sequence data for all three metabolic models under study, covering all genes in each model. For genes lacking corresponding AA sequences in the database, we use dummy sequences as placeholders.
To mitigate class imbalance between deleted and non-deleted genes, we first applied gene-wise balancing: for genes with both deletion states, an equal number of samples was randomly drawn from each state; otherwise, all available instances were retained. Next, global balancing was performed by downsampling the majority class to match the minority, resulting in a fully balanced training set. Due to metabolite number constraints, the final sample sizes approximate, rather than exactly match, the intended ratios. The details of the constructed datasets for the three metabolic models are summarized in Table VI.
TABLE VI: Summary of constructed datasets for three constraint-based metabolic models.
@@ IV-C Baseline Method: DNN
In the first set of experiments, i.e., performance evaluation of gene deletion strategy prediction methods, we employed a deep learning–based baseline: a Deep Neural Network (DNN) [DNN] to predict gene deletion status. DNNs are widely used as standard baselines in supervised learning tasks because they can capture non-linear relationships in high-dimensional data, even with relatively simple architectures.
@@
The concatenated embedding $\mathbf{X}$} is passed through two fully connected layers with ReLU activations, followed by a sigmoid output layer that produces the prediction score:
$\left\{\begin{aligned} \mathbf{h}_{1}&=\mathrm{ReLU}(\mathbf{W}_{1}\mathbf{X}+\mathbf{b}_{1}),\\ \hat{y}&=\sigma(\mathbf{W}_{2}\mathbf{h}_{1}+\mathbf{b}_{2}),\end{aligned}\right$. (26)
where $\hat{y}=P(y=1\mid\mathbf{T}_{g},\mathbf{T}_{m};\theta)$, $y\in\{0,1\}$ is the ground-truth label (deleted or non-deleted), and $\theta=\{\mathbf{W}_{1},\mathbf{W}_{2},\mathbf{b}_{1},\mathbf{b}_{2}\}$ denotes the set of learnable parameters. The DNN is trained with the same experimental setup as our proposed framework, using the binary cross-entropy loss defined in Equation 23.
@@ IV-D Baseline Method: DeepGdel
Besides the DNN baseline, we also employ DeepGdel [yang2025deepgdel] as an advanced baseline for gene deletion strategy prediction. DeepGdel extends simple DNN approaches by incorporating sequential learning through LSTM-based autoencoders for both gene and metabolite representations, allowing the model to capture long-range dependencies within the input sequences.
Specifically, DeepGdel learns metabolite and gene representations from SMILES and AA sequences using an encoder-decoder architecture:
$\mathbf{Z}_{\text{meta}}$ $=\text{LSTM-Encoder}(\mathbf{T}_{m}),$
$\mathbf{Z}_{\text{gene}}$ $=\text{LSTM-Encoder}(\mathbf{T}_{g}),$
where $\mathbf{Z}_{\text{meta}}$ and $\mathbf{Z}_{\text{gene}}$ are the learned latent representations for metabolites and genes, respectively. These representations are integrated via element-wise multiplication:
$\mathbf{Z}=\mathbf{Z}_{\text{meta}}\odot\mathbf{Z}_{\text{gene}},$
and passed through a fully connected layer with sigmoid activation for the final prediction:
$\hat{y}=\sigma(\mathbf{W}\mathbf{Z}+\mathbf{b}),$
where $\hat{y}=P(y=1\mid\mathbf{T}_{g},\mathbf{T}_{m};\theta)$ denotes the probability of gene deletion.
@@
IV-E Evaluation Metrics
We evaluate model performance using four metrics: overall accuracy, macro-averaged precision, recall, and F1 score. Definitions of these metrics are provided below.
Overall Accuracy: Accuracy measures the proportion of correctly classified instances out of the total instances. In our experiments, overall accuracy represents the proportion of genes correctly classified as deleted or non-deleted across all metabolites. It is given by:
$\text{Overall Accuracy}=\frac{\sumop\displaylimits_{m=1}^{\left\lvert\textit{M}\right\rvert}\sumop\displaylimits_{i=0}^{1}(TP_{m_{i}}+TN_{m_{i}})}{\left\lvert\textit{M}\right\rvert\times{\left\lvert\textit{G}\right\rvert}}$
Macro-Averaged Precision: Precision measures the proportion of true positive predictions out of all positive predictions for each class. The macro-averaged precision is the average of precision values across both classes (deleted and non-deleted genes) and is calculated as:
$\text{Macro-Averaged Precision}=\frac{1}{\left\lvert\textit{M}\right\rvert}\sumop\displaylimits_{m=1}^{\left\lvert\textit{M}\right\rvert}\frac{1}{2}\sumop\displaylimits_{i=0}^{1}\frac{TP_{m_{i}}}{TP_{m_{i}}+FP_{m_{i}}}$
Macro-Averaged Recall: Recall (or sensitivity) measures the proportion of true positives out of all actual positives for each class. The macro-averaged recall is the average of recall values across both classes and is given by:
$\text{Macro-Averaged Recall}=\frac{1}{\left\lvert\textit{M}\right\rvert}\sumop\displaylimits_{m=1}^{\left\lvert\textit{M}\right\rvert}\frac{1}{2}\sumop\displaylimits_{i=0}^{1}\frac{TP_{m_{i}}}{TP_{m_{i}}+FN_{m_{i}}}$
Macro-Averaged F1 Score: The F1 score is the harmonic mean of precision and recall for each class. The macro-averaged F1 score is the average of F1 scores across both classes and is given by:
$\text{Macro-Averaged F1 Score}=\frac{1}{\left\lvert\textit{M}\right\rvert}\sumop\displaylimits_{m=1}^{\left\lvert\textit{M}\right\rvert}\frac{1}{2}\sumop\displaylimits_{i=0}^{1}2\cdot\frac{\text{Prec.}_{m_{i}}\cdot\text{Recall}_{m_{i}}}{\text{Prec.}_{m_{i}}+\text{Recall}_{m_{i}}}$
Macro-Averaged AUC: The Area Under the Receiver Operating Characteristic Curve (AUC) quantifies the ability of a model to discriminate between the two classes. For each metabolite $m$, the AUC is computed based on the predicted scores and true labels. It is given by:
$\text{AUC}_{m}=\frac{1}{\left\lvert P_{m}\right\rvert\left\lvert N_{m}\right\rvert}\sumop\displaylimits_{p\in P_{m}}\sumop\displaylimits_{n\in N_{m}}\left[\mathbb{I}(s_{p}>s_{n})+\frac{1}{2}\mathbb{I}(s_{p}=s_{n})\right],$
where $P_{m}$ and $N_{m}$ are the sets of indices of positive (deleted) and negative (non-deleted) genes for metabolite $m$, $s_{p}$ and $s_{n}$ are the predicted scores (probabilities) for gene $p$ and $n$, respectively, $\mathbb{I}(\cdot)$ is the indicator function that returns 1 if the condition holds, and 0 otherwise. The macro-averaged AUC is the average of the AUC scores across all metabolites:
[…]
TABLE VII: Comparison of gene deletion strategy prediction performance between the proposed GraphGDel framework and two baseline methods across three metabolic models. “M-Avg.” denotes macro-averaged results. Best-performing values are shown in bold.
@@ IV-E Evaluation Metrics
DNN DeepGdel Proposed DNN DeepGdel Proposed DNN DeepGdel Proposed
Overall Accuracy (%) 63.31 $\pm$ 2.70 71.18 $\pm$ 0.44 77.35 $\pm$ 1.45 69.28 $\pm$ 1.58 80.58 $\pm$ 0.44 85.54 $\pm$ 1.04 60.04 $\pm$ 1.04 67.91 $\pm$ 0.34 73.22 $\pm$ 1.02
M-Avg. Precision (%) 64.99 $\pm$ 2.57 72.44 $\pm$ 0.56 75.99 $\pm$ 1.33 72.07 $\pm$ 1.44 77.51 $\pm$ 0.42 83.19 $\pm$ 1.07 64.60 $\pm$ 1.13 68.90 $\pm$ 0.35 74.67 $\pm$ 1.09
M-Avg. Recall (%) 64.13 $\pm$ 2.20 72.87 $\pm$ 0.54 75.38 $\pm$ 1.37 68.52 $\pm$ 1.36 77.85 $\pm$ 0.44 82.91 $\pm$ 1.12 61.80 $\pm$ 1.11 64.40 $\pm$ 0.35 72.51 $\pm$ 1.06
M-Avg. F1 Score (%) 59.59 $\pm$ 2.36 70.85 $\pm$ 0.65 74.45 $\pm$ 1.24 64.84 $\pm$ 1.56 76.48 $\pm$ 0.45 81.87 $\pm$ 1.25 58.69 $\pm$ 1.17 63.22 $\pm$ 0.34 72.31 $\pm$ 1.03
M-Avg. AUC (%) 62.66 $\pm$ 2.89 71.16 $\pm$ 0.85 76.56 $\pm$ 1.26 66.50 $\pm$ 1.47 80.24 $\pm$ 0.78 84.24 $\pm$ 1.12 60.30 $\pm$ 1.14 65.78 $\pm$ 0.54 73.17 $\pm$ 1.04
TABLE VIII: Ablation study of the proposed framework comparing two variants: Ge-SMILES (using only gene and SMILES-based metabolite features) and Ge-Graph (using only gene and graph-based metabolite features). Results are evaluated across the three metabolic models under study. “M-Avg.” denotes macro-averaged results. Best-performing values are shown in bold.
@@
TABLE IX: Performance comparison of the proposed framework using different graph representations ($\mathcal{G}_{0}$ and $\mathcal{G}_{1}$) across three metabolic models. “M-Avg.” denotes macro-averaged results. The best-performing values are shown in bold.
@@
IV-F Results and Performance Comparison
The results of the three computational experiments, namely (1) Performance Evaluation, (2) Module Ablation, and (3) Graph Representation Evaluation, are summarized in Tables VII, VIII, and IX, respectively.
Performance Evaluation: Across all metabolic models, GraphGDel consistently outperformed the baseline methods in overall accuracy, macro-averaged precision, recall, F1 score, and AUC. Specifically, the overall accuracy of GraphGDel was 77.35%, 85.54%, and 73.22% for e_coli_core, iMM904, and iML1515, respectively, representing substantial improvements over DNN (63.31%, 69.28%, 60.04%) and DeepGdel (71.18%, 80.58%, 67.91%). This demonstrates the robustness of GraphGDel across models of varying scales.

### v7

@@ III-A Constructing Graph Representations of Metabolic Models
As a result, the proposed graph construction pipeline produces a representation that preserves key functional structures while minimizing the confounding influence of topological hubs. The pseudocode of the proposed graph construction pipeline is presented below.
@@
Input: $\mathcal{C}=\{\textit{M},\textit{R},\textit{S},\textit{G},\textit{F}\}$, Currency Metabolite Set $\mathcal{U}$.
@@ III-C1 Metabolite Representation Learning
The Meta-M module is designed to transform raw character strings input, i.e., the SMILES (Simplified Molecular Input Line Entry System) sequences of metabolites into dense, informative feature vectors [39]. To effectively capture the complex patterns within these chemical representations, we employ a sequence-to-sequence autoencoder architecture based on Long Short-Term Memory (LSTM) networks [7, 30]. We selected LSTM as the default model because our goal is to propose a general learning framework for this task and to validate its effectiveness. The LSTM meets these requirements: it is straightforward to implement and provides a strong baseline for future comparisons with more complex replacements, such as Transformer-based models [36].
@@ III-D2 Data Partitioning Strategy
Model performance was evaluated using repeated stratified 10-fold cross-validation. The dataset was randomly partitioned into 10 folds while preserving class proportions. In each fold, approximately 90% of the data were used for training and 10% for validation.
This procedure was repeated five times with different random splits, resulting in a total of $5\times 10$ cross-validation runs. All reported performance metrics correspond to these repeated cross-validation experiments and are presented as the mean $\pm$ standard deviation across folds.
@@ III-D3 Hyperparameters
Hyperparameters were tuned via grid search within the repeated 10-fold stratified cross-validation framework (5 repeats with different random seeds) on the training/validation data. The configuration achieving the highest average validation performance across folds was selected as the final setting. The detailed search space and results are summarized in the Appendix.
@@ IV-A Overview
  • e_coli_core [19] is a simplified metabolic model of Escherichia coli (E. coli) that captures only the essential pathways necessary for core metabolic functions. It is widely used for foundational studies of bacterial metabolism and gene interactions.
  • iMM904 [14] is a more comprehensive model of Saccharomyces cerevisiae (yeast), incorporating a larger set of reactions and metabolites. It is commonly used for analyzing complex eukaryotic metabolic networks.
  • iML1515 [15] is one of the most detailed metabolic models of E. coli, covering a vast array of metabolic reactions and pathways. It provides a more intricate network compared to e_coli_core and is widely used as a genome-scale model for simulations.
@@
The detailed breakdown of genes, metabolites, and reactions in the studied models is summarized in Table III.
TABLE III: Summary of constraint-based metabolic models used in the computational experiments.
@@ IV-B1 Gene Deletion Strategy Data
Data Source. The gene deletion strategy data used in this study are obtained from the MetNetComp database [33]. MetNetComp curates systematically simulated results derived from constraint-based models and computes growth-coupled gene deletion strategies that are minimal or maximal with respect to the number of deleted genes.
@@ IV-B2 SMILES Data
Data Source. The SMILES data used in this study is obtained from the MetaNetX database [16]. MetaNetX is an online platform for accessing, analyzing, and manipulating genome-scale metabolic networks and biochemical pathways. It curates information from public repositories to integrate data on metabolites and related reactions across many species. For all three metabolic models under study, we download SMILES data for metabolites corresponding to those with available gene deletion strategy data. For metabolites lacking corresponding SMILES entries in the database, we use dummy SMILES sequences as placeholders.
@@ IV-B3 Amino Acid Sequence Data
Data Source. The AA sequence data used in this study is obtained from the KEGG database [10]. KEGG provides comprehensive information on protein-coding genes, their products, and associated metabolic pathways. The AA sequence data in KEGG is curated from publicly available genomic resources and is continuously updated to ensure accuracy and relevance. We download AA sequence data for all three metabolic models under study, covering all genes in each model. For genes lacking corresponding AA sequences in the database, we use dummy sequences as placeholders.
To mitigate class imbalance between deleted and non-deleted genes, we first applied gene-wise balancing: for genes with both deletion states, an equal number of samples was randomly drawn from each state; otherwise, all available instances were retained. Next, global balancing was performed by downsampling the majority class to match the minority, resulting in a fully balanced training set. Due to metabolite number constraints, the final sample sizes approximate, rather than exactly match, the intended ratios. The details of the constructed datasets for the three metabolic models are summarized in Table IV.
TABLE IV: Summary of constructed datasets for three constraint-based metabolic models.
@@ IV-C Baseline Method: DNN
In the first set of experiments, i.e., performance evaluation of gene deletion strategy prediction methods, we employed a deep learning–based baseline: a Deep Neural Network (DNN) [22] to predict gene deletion status. DNNs are widely used as standard baselines in supervised learning tasks because they can capture non-linear relationships in high-dimensional data, even with relatively simple architectures.
@@
The concatenated embedding $\mathbf{X}$} is passed through two fully connected layers with ReLU activations, followed by a sigmoid output layer that produces the prediction score. The DNN is trained with the same experimental setup as our proposed framework, using the binary cross-entropy loss defined in Equation 23.
@@ IV-D Baseline Method: DeepGdel
Besides the DNN baseline, we also employ DeepGdel [42] as an advanced baseline for gene deletion strategy prediction. DeepGdel extends simple DNN approaches by incorporating sequential learning through LSTM-based autoencoders for both gene and metabolite representations, with details introduced in the Appendix.
@@
IV-E Baseline Method: NMA (Neighborhood Mean Aggregation)
To evaluate the benefit of learned graph representations over a simple topology-aware scheme, we include an NMA (Neighborhood Mean Aggregation) baseline. NMA uses the same Gene-M and Meta-M modules as DeepGdel and the same metabolite graph structure as in the proposed framework (Section III-A). For prediction, instead of using the metabolite’s own Meta-M embedding, NMA aggregates Meta-M outputs over the graph neighborhood: for metabolite node $i$ with neighbor set $\mathcal{N}(i)$,
$\mathbf{z}_{i}^{\mathrm{nma}}=\frac{1}{|\mathcal{N}(i)|}\sumop\displaylimits_{j\in\mathcal{N}(i)}\mathbf{z}_{j}^{\mathrm{meta}},$
where $\mathbf{z}_{j}^{\mathrm{meta}}$ is the Meta-M embedding of metabolite $j$; if $|\mathcal{N}(i)|=0$, $\mathbf{z}_{i}^{\mathrm{nma}}=\mathbf{z}_{i}^{\mathrm{meta}}$. The (gene, metabolite) prediction is then
$\hat{y}=\sigma\bigl(\mathbf{w}^{\top}(\mathbf{h}_{g}\odot\mathbf{z}_{i}^{\mathrm{nma}})+b\bigr),$
with $\mathbf{h}_{g}$ the Gene-M embedding and $\odot$ the element-wise product. NMA is trained using the same protocol as the other baselines and serves as a topology-aware, non-learned graph aggregation baseline for comparison with our graph-learning method.
IV-F Evaluation Metrics
To ensure a fair comparison, the same data partitioning strategy and hyperparameter tuning protocol (grid search with 10-fold stratified cross-validation, repeated five times with different random seeds) were applied to the proposed framework and to all baseline methods (DNN, DeepGdel, and NMA) in the performance evaluation experiments; the best configuration for each method was selected based on validation performance. We evaluate model performance using four metrics: Overall Accuracy, Macro-averaged Precision, Recall, F1 Score, and AUC. The detailed definitions and formulations of these metrics can be found in the Appendix.
TABLE V: Comparison of gene deletion strategy prediction performance between the proposed GraphGDel framework and baseline methods (DNN, DeepGdel, NMA) across three metabolic models. “M-Avg.” denotes macro-averaged results. Best-performing values are shown in bold.
@@ IV-F Evaluation Metrics
DNN DeepGdel NMA Proposed DNN DeepGdel NMA Proposed DNN DeepGdel NMA Proposed
Overall Accuracy (%) 63.31 $\pm$ 2.70 71.18 $\pm$ 0.44 72.25 $\pm$ 0.52 77.35 $\pm$ 1.45 69.28 $\pm$ 1.58 80.58 $\pm$ 0.44 81.18 $\pm$ 0.50 85.54 $\pm$ 1.04 60.04 $\pm$ 1.04 67.91 $\pm$ 0.34 68.52 $\pm$ 0.38 73.22 $\pm$ 1.02
M-Avg. Precision (%) 64.99 $\pm$ 2.57 72.44 $\pm$ 0.56 73.12 $\pm$ 0.58 75.99 $\pm$ 1.33 72.07 $\pm$ 1.44 77.51 $\pm$ 0.42 78.38 $\pm$ 0.45 83.19 $\pm$ 1.07 64.60 $\pm$ 1.13 68.90 $\pm$ 0.35 69.48 $\pm$ 0.38 74.67 $\pm$ 1.09
M-Avg. Recall (%) 64.13 $\pm$ 2.20 72.87 $\pm$ 0.54 73.42 $\pm$ 0.56 75.38 $\pm$ 1.37 68.52 $\pm$ 1.36 77.85 $\pm$ 0.44 78.52 $\pm$ 0.48 82.91 $\pm$ 1.12 61.80 $\pm$ 1.11 64.40 $\pm$ 0.35 65.18 $\pm$ 0.38 72.51 $\pm$ 1.06
M-Avg. F1 Score (%) 59.59 $\pm$ 2.36 70.85 $\pm$ 0.65 71.52 $\pm$ 0.68 74.45 $\pm$ 1.24 64.84 $\pm$ 1.56 76.48 $\pm$ 0.45 77.28 $\pm$ 0.48 81.87 $\pm$ 1.25 58.69 $\pm$ 1.17 63.22 $\pm$ 0.34 64.08 $\pm$ 0.36 72.31 $\pm$ 1.03
M-Avg. AUC (%) 62.66 $\pm$ 2.89 71.16 $\pm$ 0.85 72.08 $\pm$ 0.88 76.56 $\pm$ 1.26 66.50 $\pm$ 1.47 80.24 $\pm$ 0.78 81.02 $\pm$ 0.82 84.24 $\pm$ 1.12 60.30 $\pm$ 1.14 65.78 $\pm$ 0.54 66.52 $\pm$ 0.56 73.17 $\pm$ 1.04
To provide additional insight into model behavior and error distribution, we report confusion matrices for the proposed method and the baseline methods (DNN, DeepGdel, and NMA) on the three metabolic models in Table VI. Table VI reports row-normalized proportions for each true label. The positive class corresponds to “non-deleted” (gene not deleted) and the negative class to “deleted.”
TABLE VI: Row-normalized confusion matrices of all methods across three metabolic models. Rows correspond to true labels and columns to predicted labels. Values are row-wise proportions.
Method True label e_coli_core iMM904 iML1515
Pred. Del. Pred. N-Del. Pred. Del. Pred. N-Del. Pred. Del. Pred. N-Del.
DNN Del. 0.69 0.31 0.71 0.29 0.55 0.45
N-Del. 0.42 0.58 0.34 0.66 0.31 0.69
DeepGdel Del. 0.76 0.24 0.88 0.12 0.78 0.22
N-Del. 0.34 0.66 0.33 0.67 0.49 0.51
NMA Del. 0.79 0.21 0.90 0.10 0.77 0.23
N-Del. 0.33 0.67 0.30 0.70 0.40 0.60
Proposed Del. 0.83 0.17 0.93 0.07 0.75 0.25
N-Del. 0.32 0.68 0.27 0.73 0.30 0.70
TABLE VII: Ablation study of the proposed framework comparing two variants: Ge-SMILES (using only gene and SMILES-based metabolite features) and Ge-Graph (using only gene and graph-based metabolite features). Results are evaluated across the three metabolic models under study. “M-Avg.” denotes macro-averaged results. Best-performing values are shown in bold.
@@
TABLE VIII: Performance comparison of the proposed framework using different graph representations ($\mathcal{G}_{0}$ and $\mathcal{G}_{1}$) across three metabolic models. “M-Avg.” denotes macro-averaged results. The best-performing values are shown in bold.
@@
IV-G Results and Performance Comparison
The results of the three computational experiments, namely (1) Performance Evaluation, (2) Module Ablation, and (3) Graph Representation Evaluation, are summarized in Tables V, VII, and VIII, respectively. To assess whether the reported differences are statistically significant, we performed paired $t$-tests on the cross-validation fold-level results (50 paired values per comparison, obtained from 5 repeats $\times$ 10 folds), comparing the proposed method against each baseline. All comparisons yielded $p<0.05$ across every metric and model (Tables V, VII, and VIII).
Performance Evaluation: Across all metabolic models, GraphGDel consistently outperformed the baseline methods in overall accuracy, macro-averaged precision, recall, F1 score, and AUC. Specifically, the overall accuracy of GraphGDel was 77.35%, 85.54%, and 73.22% for e_coli_core, iMM904, and iML1515, respectively, representing substantial improvements over DNN (63.31%, 69.28%, 60.04%), DeepGdel (71.18%, 80.58%, 67.91%), and NMA (72.25%, 81.18%, 68.52%). This demonstrates the robustness of GraphGDel across models of varying scales.